OpenStudy (anonymous):

Find the equation of the line that contains the point (3, 5) and is parallel to the line 2x + y = -4.

OpenStudy (anonymous):

plug the point (3,5) into the equation 2x + y = k and solve for k your answer will be 2x + y = whatever you came up with for k

OpenStudy (anonymous):

ok but then what happens to the neg4

OpenStudy (anonymous):

just forget it in this particular type of problem

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

which I get 11

OpenStudy (anonymous):

2x+y describes the slope... -4 is the y intercept so 2x + y = -4 is parrallel and below 2x + y = 0 which is parrallel and below 2x + y = 11

OpenStudy (anonymous):

exactly right... good job

OpenStudy (anonymous):

thank you